Please help me solve this word problems:Mr.Avila is four times as old as his son. In 16 years he will be only twice as old .What are their ages?
Answer by Cromlix(4381)

Mr.Avila is four times as old as his son.
A = 4S
In 16 years he will be only twice as old
A + 16 = 2(S + 16)
Replace A with 4S (from 1st equation)
4S + 16 = 2(S + 16)
4S + 16 = 2S + 32
Collect like terms.
4S - 2S = 32 - 16
2S = 16
S = 8
Therefore, Mr Avila is 32 yrs old
and his son is 8 years old.
Hope this helps.
:-)